# Feature-Based Interpretable Optimization

### v1

@@ 4.1 Learning Heuristic
In this heuristic, we first generate training data by solving the underlying optimization problem for each of the $N$ available candidate scenarios. We then extract the meta-solutions corresponding to these $N$ optimal solutions. Of these at most $N$ different meta-solutions, we choose the $K$ best, which then are used to find a classification tree. Algorithm 1 describes the heuristic in detail for minimization problems.
$j\in[N]$
[…]
4:$\ell_{j}\leftarrow$ index $i\in[N]\ s.t.\ \mu_{ij}=1$EndFor
[…]
@@
In this heuristic, we first solve the nominal problem for each scenario $j$ (line 1) to find solution $\bm{x}^{j}$. Extracting its features, there is a corresponding meta-solution. We then evaluate this meta-solution for each scenario to find its objective value (line 2). If a problem is infeasible, we set $C^{*}_{ij}=\infty$. Note that we may obtain the same solution $\bm{x}^{j}$ for multiple scenarios $j\in[N]$, in which case some of the evaluation loops in line 1-2 can be skipped.
@@
Having obtained a choice of $K$ meta-solutions this way, we still need to construct a decision tree. For this purpose, we mark the best meta-solution per scenario (line 4) and train a standard classification tree with limited depth, e.g., using scikit-learn’s implementation of the CART algorithm.
@@ 4.2 M2M Heuristic
A second heuristic that we use is based on the approach presented in the paper [GH23]. Their presented MIP formulation and greedy heuristic aim to create a decision tree and identify solutions (not meta-solutions) for the scenarios assigned to its leaves. The micro-to-macro (M2M) heuristic, first uses the MIP or heuristic from the original framework and replaces the solutions at the leaves with the respective meta-solution. For all experiments presented in this paper, their MIP formulation was used to obtain the decision trees.
@@ 4.3 Best single solutions
In this section, we consider the impact of using a single solution to optimize performance over all the scenarios. We consider the case of the best meta-solution and the best micro-solution.
@@
$\min_{\phi_{S}\in\Phi_{S}}\mu_{\bm{p}}\left(C(\phi_{I}^{1},\phi_{S}),\ldots,C(\phi_{I}^{N},\phi_{S})\right).$
The resulting optimization problem avoids finding an optimization rule and only provides a meta-solution to be applied in every scenario. These meta-solutions can be obtained by solving the MIP formulations 2 and 3 for $K=1$ which allows removing the variables and constraints regarding the tree structure as well as index $k$.
@@
$\min_{\bm{x}\in{\mathcal{X}}}\mu_{\bm{p}}\left(c(\phi_{I}^{1},\bm{x}),\ldots,c(\phi_{I}^{N},\bm{x})\right).$
Note that this is only applicable if ${\mathcal{X}}={\mathcal{X}}(\phi_{I}^{i})$ for all scenarios $i\in[N]$, i.e., the optimization domain is the same for all scenarios.
@@ 5.1 Setting
We denote the different models as NAMEx where NAME indicates the method and x indicates the number of leaves of the decision trees. For both sets of problems the following methods were used: the MIPs presented in Section 3.2 and 3.3 (MIPx), the learning heuristic from Section 4.1 (LHx) and the M2M heuristic from Section 4.2 (M2Mx) for generating trees with meta-solutions. For the use of M2Mx and as a benchmark the MIP formulation from [GH23] (MICROx) was also implemented. To ensure interpretability, we restricted ourselves to trees with two and four leaves. For comparisons, we also compute the single best meta-solution (referred to as META01) and the single best micro-solution (referred to as MICRO1) as described in Section 4.3. Since here all scenarios are mapped to the same (meta-)solution, both of them can be considered as being the most interpretable of our methods. In contrast, solving every scenario to optimality (OPT) using a black-box solver instead of any mapping represents a non-interpretable method.
In addition to the experiments using all methods, we also conducted experiments using larger instances. In the latter, the MIPx methods were not used due to their intractability. To generate trees with two leaves, we used the formulation for symmetric trees (1), for trees with four leaves we used a formulation for asymmetric trees since its use has proven to be slightly faster.
For both problem types the values for the vectors $\bm{c}$, representing the costs of edges respectively items, were generated using a variation of the approach from [GH23]. Three distributions were generated, which we used as base scenarios. The costs of each edge or item were uniformly sampled from one of these distributions, which was chosen randomly with equal probabilities. To generate test sets, we sampled scenarios in the same way. Each test set contains 100 scenarios. More information about the generation of the problem-specific parameters of the instances is included in the respective subsections.
@@
A value of one thus represents the objective that can be achieved if each instance is solved by itself to optimality, and is, therefore, an upper bound for all methods used. A higher value hence indicates even for minimization problems (i.e. in our experiments shortest path) a better solution.
To evaluate the generated optimization rules, it is necessary to identify the corresponding leaf for a given scenario and compute a feasible micro-solution based on the meta-solution found in that leaf. All our experiments are based on the assumption that the features have been selected reasonably, which allows a stakeholder to practically work with given meta-solutions. Therefore, we simulate that the user can find an optimal solution for a combination of meta-solution and instance. The resulting optimization problem of finding a micro-solution for the respective meta-solution was thus always solved to optimality.
We performed 100 runs for each data point presented in the following. To ensure reproducibility, the seeds for generating the instances were fixed. The running times presented include both the time required to solve and construct the problem. We implemented all methods and generated the data using Python version 3.11. As part of the implementation of the learning heuristic (LHx) we used the function DecisionTreeClassifier from the scikit-learn library for generating the classification trees. We also used the Networkx library for storing and handling graph structures. All (M)IP formulations were solved using gurobipy and Gurobi version 11. Its internal time limit was set to 900 seconds. The experiments presented in 5.2 to 5.3.2 were conducted on a cluster of machines with 2.6-3.3 GHz Intel Xeon CPUs with each 4 cores and 12 GB memory. For the examples illustrated in 5.3.3 a machine with an Apple M2 Pro chip with ten cores and 16 GB of memory was used.
@@ 5.2 Knapsack Problems
In this section, we present computational experiments using our framework for the knapsack problem. We conducted experiments where the number of items $n$ (Section 5.2.1), the number of used training scenarios $N$ (Section 5.2.2) and the dimension of features $F$ (Section 5.2.3) were varied.
All experiments were conducted using artificially generated data. The weights $w_{i}$ were uniformly sampled from the interval $[0.1,10]$. For our experiments we considered a fixed budget $C$ which was set to $\frac{1}{2}\sum_{i\in[n]}w_{i}$. For formulating the meta-solutions, the items were grouped as evenly as possible in $F$ sets. As described in Section 3.2 the budgets $C^{k}_{f}$ which are chosen as part of the meta-solutions do not necessarily sum up to the budget given as part of the instance $C$. Therefore, those were scaled before evaluation such that $\sum_{f\in[F]}C^{k}_{f}=C$ for all $k\in[K]$ while retaining their ratios.
5.2.1 Varying Number of Items
First, we present experiments in which the number of items was varied. For these, the number of training scenarios $N$ was set to ten, and the number of features $F$ was set to four. Figures 10(a) and 10(b) illustrate the scaled objective value on the training data in dependency of the number of items provided per instance. In Figure 10(a) the solid green and red lines are partially obscured by the blue line. It can be observed that there is a negative correlation between the number of items and the objective for all trees with two and four leaves. This trend is more distinct for trees with two leaves. Moreover, the methods that generate trees with four leaves demonstrate considerably better performance. However, it is remarkable that those from LH4 perform significantly worse than the other methods including the benchmark. This can be explained by the fact that the trees are constructed with regard to the accuracy and not the objective function value of the knapsack problems. As illustrated in Figure 10(b), the performance of LH2 and LH4 as well as the performance of M2M2 increases again as the number of items rises. Even with the performance of M2M2 increasing, for 140 and more items this method is not able to outperform META1. This means that for these instances the tree structure cannot be used beneficially. Moreover, the quality of the solutions from MICRO2 and MICRO4 continues to decline. Notably, the performance differences between MICRO2 and M2M2 are considerably larger than those between MICRO4 and M2M4.
(a) objective training/ #items.
(b) objective training/ #items.
(c) objective test/ #items.
(d) objective test/ #items.
(e) runtime/ #items.
(f) runtime/ #items.
@@ 5.2.1 Varying Number of Items
Refer to caption MIP2 Refer to caption MICRO2 Refer to caption LH2 Refer to caption M2M2
@@
Figure 10: Plots for knapsack and synthetic data, #items/ on x-axis.
Figures 10(c) and 10(d) depict the relationship between the scaled test objective and the number of items per instance. For all methods except MICRO2 and MICRO4, a positive correlation can be identified. For instances with a small number of items, the methods MIP4, MICRO4, M2M4 are overfitting. The gap between MICROx and M2Mx derived from it increases significantly with an increasing number of items. In contrast to its performance on the training data, the performance on the test data of LH2 and LH4 is highly competitive. For instances with many items, significantly better performance is achieved when using META1 than MICROx. In general, it can be concluded that the performance of our methods appears to be clearly advantageous in terms of performance on the test data.
The median computing time used is shown in Figures 10(e) and 10(f). The time required for solving via M2Mx results almost exclusively from the time used to calculate MICROx. The data points of MICROx therefore also represent the computing time of M2Mx. Even small instances for MIP4 could not be solved to optimality without hitting the time limit of 900 seconds. It can even be seen that all MIP-based methods except MICRO2 cannot be solved to optimality within the time limit. In every run, feasible solutions were found within the time limit. LHx can be solved for all configurations within seconds. For both MICRO2 and MICRO4, dips in the progressions can be observed.
@@ 5.2.2 Varying Number of Scenarios
Figures 11(e) and 11(f) show the computation time for a varying number of scenarios with $n=16$ and $F=4$. For MIP4 even instances with ten scenarios couldn’t be solved to optimality. Also, using MIP2 and MICRO4 the time limit was hit even for 16 respective 18 training scenarios. The computation times for MICRO2 show a notable dip from 14 to 23 scenarios, which can not be explained. In contrast to the MIP-based approaches, using LH2 and LH4 even for instances with 200 scenarios trees can be generated in less than two seconds.
@@
Refer to caption MIP2 Refer to caption MICRO2 Refer to caption LH2 Refer to caption M2M2
@@
Figure 11: Plots for knapsack and synthetic data, #scenarios/ on x-axis.
The relationship between the scaled objective on the training set and the number of scenarios used for training is shown in Figures 11(a) and 11(b). For all methods generating trees with two leaves, the influence on the training performance is marginal. The performance of MICRO4 and M2M4 decreases significantly when more than 60 scenarios are used, and even falls below that of MICRO2 and M2M2. This can be explained because no (proven) optimal solutions could be determined for MICRO4. It can be seen that the quality of the solutions of MICRO4 and M2M4 is strongly correlated and that of M2M4 presumably depends on MICRO4.
In Figure 11(c) it can be seen that for all methods utilized, the performance increases with the number of scenarios used for training. In contrast, Figure 11(d) illustrates that the performance of MICRO4 and M2M4 begins to decline at a certain point as the number of scenarios increases. As with the data in Fig. 11(b), this is due to the intractability of the models. It is noteworthy that LH4 clearly outperforms MIP4. In general, all developed methods outperform the benchmark model of the same size. Furthermore, it can be seen that the marginal benefit of further training scenarios decreases.
5.2.3 Varying Number of Features
For the last set of knapsack experiments we set the number of training scenarios to ten, the number of items to 16, and varied the number of features. The results are illustrated in Figures 12(a), 12(b) and 12(c). These experiments include two noteworthy special cases: Firstly, the case where the number of features is set to one. In this case, every given solution space is equal to the set of feasible solutions of the knapsack problem. Due to our assumption, that the user is able to find an optimal solution for a given solution space, the methods M2Mx, LHx and MIPx results in a scaled objective of one for the training as well as the test data. The second case includes the experiments where the number of features was set equal to the number of items. A given meta-solution therefore consists of individual budgets for every single item and represents consequentially a micro-solution for the knapsack problem. This is reflected in the results to the extent that MIPx, MICROx and M2Mx have in this case the same objective on the training data. As expected, the gap between the benchmark and our methods closes with an increasing number of features i.e., a competent user can benefit from a more rough description of the solution space which results in a higher degree of freedom. On the other hand, for an individual who is less familiar with the problem domain, an increasing degree of freedom (i.e. small value $F$) could lead to worse solutions.
(a) objective training/ #features.
(b) test objective/ #features.
(c) runtime/ #features.
Refer to caption MIP2 Refer to caption MICRO2 Refer to caption LH2 Refer to caption M2M2
Refer to caption META1
Figure 12: Plots for knapsack and synthetic data, #features/ on x-axis.
In terms of objective on the test data, it can be seen that the LHx methods clearly outperform all the other methods. It is especially remarkable, that the use of trees with only two leaves generated by LH2 nearly always results in better objective values than the use of trees with four leaves made by MP4. Furthermore, it is noticeable that all the MIP-based formulations can be solved particularly quickly for the two extreme cases described.
5.3 Shortest Path Problems
In this section, we present results generated by solving shortest-path problems. For artificial graphs, we conducted experiments where the number of scenarios $N$ (Section 5.3.1) as well as the dimension of features $F$ and the number of nodes $n\times n$ (Section 5.3.2) were varied.
Those were performed on grid graphs with $n\times n$ nodes, where the node in the southwest corner represents the source and the node in the northeast represents the sink. Every node is connected to its nearest neighbor in its east (if existing) and to the one in its north (if existing). The edges are directed from the south to the north and the west to the east. All graphs used are therefore directed acyclic graphs. The nodes are divided into $F$ equally sized quadratic districts, which are used to describe the meta-solutions. In this setting, every meta path from the source $s$ to the sink $t$ has the same length, which is dependent on $F$. When solving MIP2 and MIP4 the parameter $\Delta$ which regulates the maximum length of a meta-solution was therefore set to $12\sqrt{F}-$ 1.
In Section 5.3.3 we present extended results for the experiments conducted on the Chicago graph [CDG19] introduced in 2.2. In these experiments, we used ten scenarios as a training set and 1,000 scenarios as a test set. All scenarios consist of artificially generated edge costs representing four different traffic scenarios presented in Figure 1.
5.3.1 Varying Number of Scenarios
Figures 13(a) to 13(f) illustrate the relationship between objective/time and number of used training scenarios. The results show that the quality of the solutions on the training data essentially does not increase with more scenarios considered. The solution quality of MICRO2 even decreases. It can be observed that – as expected – the performance on the test data improves as the number of training scenarios increases. While for most methods this is only a slight improvement, the impact on MICROx is quite significant. As with the experiments considering knapsack problems, the required training time shows a dip for MICRO4 that we cannot explain. There is an otherwise clear positive correlation between the number of scenarios and the computing time. For all investigated instances the use of LHx is preferable to that of M2Mx.
(a) objective training/ #scenarios.
(b) objective training/ #scenarios.
(c) objective test/ #scenarios.
(d) objective test/ #scenarios.
(e) runtime/ #scenarios.
(f) runtime/ #scenarios.
Refer to caption MIP2 Refer to caption MICRO2 Refer to caption LH2 Refer to caption M2M2
Refer to caption META1
Figure 13: Plots for shortest path and synthetic data, #scenarios on x-axis.

### v3

@@ 4.1 Learning Heuristic
In this heuristic, we first generate training data by solving the underlying optimization problem for each of the $N$ available candidate scenarios. We then extract the meta-solutions corresponding to these $N$ optimal solutions. Of these, at most $N$ different meta-solutions, we choose the $K$ best, which are then used to find a classification tree. Algorithm 1 describes the heuristic in detail for minimization problems.
1: for all $j\in[N]$ do
2:   $\bm{x}^{j}\leftarrow\min_{\bm{x}\in\mathcal{X}(\bm{\phi}_{I}^{j})}c(\bm{\phi}_{I}^{j},\bm{x})$
3:   for all $i\in[N]$ do
4:    $C^{*}_{ij}\leftarrow\min_{\bm{x}\in\mathcal{X}(\bm{\phi}_{I}^{i}),\ \bm{\phi}_{S}(\bm{x})=\bm{\phi}_{S}(\bm{x}^{j})}c(\bm{\phi}_{I}^{i},\bm{x})$
5: $(\bm{\mu},\bm{\lambda})\leftarrow$ solve Problem (4) with $C^{*},K$
6: for all $j\in[N]$ do
7:   $\ell_{j}\leftarrow$ index $i\in[N]\ s.t.\ \mu_{ij}=1$
8: return classification tree on $\{(\bm{\phi}_{I}^{j},\ell_{j}):j\in[N]\}$
@@
More specifically, in Algorithm 1, we first solve the nominal problem for each scenario $j$ (line 2) to find solution $\bm{x}^{j}$. Extracting its features, there is a corresponding meta-solution. We then evaluate this meta-solution for each scenario to find its objective value (line 4). If a problem is infeasible, we set $C^{*}_{ij}=\infty$. Note that we may obtain the same solution $\bm{x}^{j}$ for multiple scenarios $j\in[N]$, in which case some of the evaluation loops in line 3-4 can be skipped.
@@
Having obtained a choice of $K$ meta-solutions this way, we still need to construct a decision tree. For this purpose, we identify the best meta-solution per scenario (line 7) and train a standard classification tree with limited depth, e.g., using scikit-learn’s implementation of the CART algorithm.
@@ 4.2 M2M Heuristic
A second heuristic that we use is based on the approach presented in Goerigk and Hartisch (2023), where a MIP formulation and a greedy heuristic were presented with the aim of creating a decision tree and identifying solutions (not meta-solutions) for the scenarios assigned to its leaves. The micro-to-macro (M2M) heuristic first uses the MIP or the heuristic from the solution-based approach to find a decision tree that can select solutions based on scenarios. It then replaces the solutions at the leaves with the respective meta-solution. For all experiments presented in this paper, the MIP formulation was used to obtain the decision trees.
@@ 4.3 Best Single Solutions
In this section, we evaluate the consequences of using only a single solution to optimize performance over all the scenarios. We consider the case of both the best meta-solution and the best micro-solution.
@@
$\min_{\bm{\phi}_{S}\in\Phi_{S}}\mu_{\bm{p}}\left(C(\bm{\phi}_{I}^{1},\bm{\phi}_{S}),\ldots,C(\bm{\phi}_{I}^{N},\bm{\phi}_{S})\right).$
The resulting optimization problem avoids finding an optimization rule and only provides one meta-solution to be applied in every scenario. Such a meta-solution can be obtained by solving the MIP formulations 2 and 3 for $K=1$, which allows removing the variables and constraints regarding the tree structure as well as index $k$.
@@
$\min_{\bm{x}\in{\mathcal{X}}}\mu_{\bm{p}}\left(c(\bm{\phi}_{I}^{1},\bm{x}),\ldots,c(\bm{\phi}_{I}^{N},\bm{x})\right).$
Note that this is only applicable if $\cap_{i\in[N]}{\mathcal{X}}(\bm{\phi}_{I}^{i})\neq\emptyset$, i.e., the intersection of the optimization domains of all scenarios is nonempty.
@@ 5.1 Setting
The methods used are presented in Table 1. To ensure interpretability, we restricted ourselves to trees with two and four leaves. In addition to the tree-based methods, we also compute the single best meta-solution (referred to as META1) and the single best micro-solution (referred to as MICRO1) as described in Section 4.3. Since here all scenarios are mapped to the same (meta-)solution, both of them can be considered as being the most interpretable of our methods. In contrast, solving every scenario to optimality (OPT) using a black-box solver instead of any mapping represents a non-interpretable method.
Method name Description Reference
MIPx MIP formulation Section 3.2 & 3.3
LHx Learning heuristic Section 4.1
M2Mx Micro-to-macro heuristic Section 4.2
MICROx Solution-based approach Goerigk and Hartisch (2023)
META1 Single best meta-solution Section 4.3
MICRO1 Single best micro-solution Section 4.3
OPT Optimal solution for each scenario
Table 1: List of methods evaluated. We denote the different models as NAMEx where NAME indicates the method and x indicates the number of leaves of the decision trees.
In addition to the experiments using all methods, we also conducted experiments using larger instances. In the latter, the MIPx methods were not used due to their intractability. To generate trees with two leaves, we used the formulation for symmetric trees (1), and for trees with four leaves, we used a formulation for asymmetric trees since its use has proven to be slightly faster.
For both problem types, the values for the vectors $\bm{c}$, representing the costs of edges respectively items, were generated using a variation of the approach from Goerigk and Hartisch (2023). Three distributions were generated, which we used as base scenarios. The costs of each edge or item were uniformly sampled from one of these distributions, which was chosen randomly with equal probabilities. To generate test sets, we sampled scenarios in the same way. Each test set contains 100 scenarios. More information about the generation of the problem-specific parameters of the instances is included in the respective subsections.
@@
A value of one thus represents the objective that can be achieved if each instance is solved by itself to optimality, and is, therefore, an upper bound for all methods used. A higher value thus indicates, even for minimization problems (i.e., in our experiments shortest path), a better solution. The gap between the value of the scaled objective and $1.0$ is an approximation of the price paid to obtain an interpretable solution instead of the true optimal solution and thus reflects the cost of interpretability. Note that the cost of interpretability arises not only from using a surrogate model but also from factors such as the quality of the implemented solution within a meta-solution and the lack of training data.
To evaluate the generated optimization rules, it is necessary to identify the corresponding leaf for a given scenario and compute a feasible micro-solution based on the meta-solution found in that leaf. All our experiments are based on the assumption that the features have been selected reasonably, which allows a stakeholder to practically work with the given meta-solutions. Therefore, we simulate that the user can find an optimal solution for a combination of meta-solution and instance. The resulting optimization problem of finding a micro-solution for the respective meta-solution was thus always solved to optimality.
We performed 100 runs for each data point presented in the following. To ensure reproducibility, the seeds for generating the instances were fixed. The running times presented include both the time required to solve and construct the problem. We implemented all methods and generated the data using Python version 3.11. As part of the implementation of the learning heuristic (LHx), we used the function DecisionTreeClassifier from the scikit-learn library (Pedregosa et al., 2011) for generating the classification trees. We also used the Networkx library (Hagberg et al., 2008) for storing and handling graph structures. All (M)IP formulations were solved using gurobipy and Gurobi version 11 (Gurobi Optimization, LLC, 2024). Its internal time limit was set to 900 seconds. The experiments presented in Sections 5.2 to 5.3.2 were conducted on a cluster of machines with 2.6-3.3 GHz Intel Xeon CPUs with 4 cores and 12 GB of memory. For the examples illustrated in Section 5.3.3, a machine with an Apple M2 Pro chip with ten cores and 16 GB of memory was used. The code and the generated data are available on GitHub.
@@ 5.2 Knapsack Problems
In this section, we present computational experiments using our framework for the knapsack problem. We place particular focus on experiments where the dimension of solution features $F_{S}$ was varied. The results of further extensive experiments, where the number of items $n$ and the number of used training scenarios $N$ were varied can be found in the Electronic Companion.
All experiments were conducted using artificially generated data. The weights $w_{i}$ were uniformly sampled from the interval $[0.1,10]$. For our experiments we considered a fixed budget $C$ which was set to $\frac{1}{2}\sum_{i\in[n]}w_{i}$. For formulating the meta-solutions, the items were grouped as evenly as possible in $F_{S}$ sets. As described in Section 3.2, the budgets $C^{k}_{f}$, which are chosen as part of a meta-solution, do not necessarily sum up to the budget given as part of the instance $C$. Therefore, those were scaled before evaluation such that $\sum_{f\in[F_{S}]}C^{k}_{f}=C$ for all $k\in[K]$ while retaining their ratios.
For the illustrated experiments, we set the number of training scenarios to ten, the number of items to 16, and varied the number of features. The results are presented in Figures 8(a), 8(b) and 8(c). These experiments include two noteworthy special cases: Firstly, the case where the number of features is set to one. In this case, every given solution space is equal to the set of feasible solutions of the knapsack problem. Due to our assumption that the user can find an optimal solution for a given solution space, the methods M2Mx, LHx and MIPx result in a normalized objective between zero and one for the training as well as the test data. The second case includes the experiments where the number of features was set equal to the number of items. A given meta-solution, therefore, consists of individual budgets for every single item and consequently represents a micro-solution for the knapsack problem. This is reflected in the results to the extent that MIPx, MICROx, and M2Mx have in this case the same objective on the training data. As expected, the gap between the benchmark and our methods closes with an increasing number of features i.e., a competent user can benefit from a rougher description of the solution space, which results in a higher degree of freedom. On the other hand, for an individual who is less familiar with the problem domain, an increasing degree of freedom (i.e., small value $F_{S}$) could lead to worse solutions.
(a) objective training/ #features.
(b) test objective/ #features.
(c) runtime/ #features.
@@
Refer to caption MIP2 Refer to caption MICRO2 Refer to caption LH2 Refer to caption $M2$
@@
Figure 8: Plots for knapsack and synthetic data, #features on x-axis. A higher scaled objective indicates a better result. The solution-based approach, which serves as a benchmark, is highlighted in red.
In terms of objective performance on the test data, it can be seen that the LHx methods clearly outperform all the other methods. It is especially remarkable that the use of trees with only two leaves generated by LH2 nearly always results in better objective values than the use of trees with four leaves made by MP4. Furthermore, it is noticeable that all the MIP-based formulations can be solved particularly quickly for the two extreme cases described.
5.3 Shortest Path Problems
In this section, we present results generated by solving shortest-path problems. For artificial graphs, we conducted experiments where the number of scenarios $N$ (Section 5.3.1) as well as the dimension of solution features $F_{S}$ and the number of nodes $n\times n$ (Section 5.3.2) were varied.
Those were performed on grid graphs with $n\times n$ nodes, where the node in the southwest corner represents the source and the node in the northeast represents the sink. Every node is connected to its nearest neighbor to its east (if one exists) and to the one in the north (if it exists). The edges are directed from the south to the north and from the west to the east. All graphs used are therefore directed acyclic graphs. The nodes are divided into $F_{S}$ equally sized square districts, which are used to describe the meta-solutions. In this setting, every meta path from the source $s$ to the sink $t$ has the same length, which is dependent on $F_{S}$. When solving MIP2 and MIP4, the parameter $\Delta$, which regulates the maximum length of a meta-solution, was therefore set to $2\sqrt{F_{S}}-1$.
In Section 5.3.3, we present extended results for the experiments conducted on Chicago’s street network (Chassein et al., 2019) introduced in 2.2. It comprises 1,308 edges and 538 nodes. In these experiments, we used ten scenarios as a training set and 1,000 scenarios as a test set. All scenarios consist of artificially generated edge costs representing four different traffic scenarios presented in Figure 2 as well as the time of the day and the day of the week, both encoded as integers.
@@ 5.3.1 Varying Number of Scenarios
Figures 9(a) to 9(f) illustrate the relationship between objective/time and the number of used training scenarios. The results show that the quality of the solutions on the training data essentially does not increase with more scenarios considered. The solution quality of MICRO2 even decreases. It can be observed that—as expected—the performance on the test data improves as the number of training scenarios increases. While for most methods this is only a slight improvement, the impact on MICROx is quite significant. As with the experiments considering knapsack problems, the required training time shows a dip for MICRO4 that we cannot explain. There is an otherwise clear positive correlation between the number of scenarios and the computing time. For all investigated instances, the use of LHx is preferable to that of M2Mx.
@@
Refer to caption MIP2 Refer to caption MICRO2 Refer to caption LH2 Refer to caption $M2$
@@
Figure 9: Plots for shortest path and synthetic data, #scenarios on x-axis. A higher scaled objective indicates a better result. The solution-based approach, which serves as a benchmark, is highlighted in red.